### v2

# Distribution-aware Knowledge Unification and Association for Non-exemplar Lifelong Person Re-identification

Abstract
Lifelong person re-identification (LReID) encounters a key challenge: balancing the preservation of old knowledge with adaptation to new information. Existing LReID methods typically employ knowledge distillation to enforce representation alignment. However, these approaches ignore two crucial aspects: specific distribution awareness and cross-domain unified knowledge learning, both of which are essential for addressing this challenge. To overcome these limitations, we propose a novel distribution-aware knowledge unification and association (DKUA) framework where domain-style modeling is performed for each instance to propagate domain-specific representations, enhancing anti-forgetting and generalization capacity. Specifically, we design a distribution-aware model to transfer instance-level representations of the current domain into the domain-specific representations with the different domain styles, preserving learned knowledge without storing old samples. Next, we propose adaptive knowledge consolidation (AKC) to dynamically generate the unified representation as a cross-domain representation center. To further mitigate forgetting, we develop a unified knowledge association (UKA) mechanism, which explores the unified representation as a bridge to explicitly model inter-domain associations, reducing inter-domain gaps. Finally, distribution-based knowledge transfer (DKT) is proposed to prevent the current domain distribution from deviating from the cross-domain distribution center, improving adaptation capacity. Experimental results demonstrate that our DKUA outperforms state-of-the-art methods on two training orders, improving anti-forgetting and generalization capacity. Our code is available at https://github.com/LiuShiBen/DKUA.
I Introduction
Figure 1: (a) Rehearsal-based approaches enforce representation alignment by storing old samples, raising data privacy concerns. (b) Rehearsal-free methods only model the style of the last domain, limiting the model’s long-term anti-forgetting capacity. (c) Our DKUA learns domain-specific styles generated by a domain-style encoder (DSE) including multiple transfer modules (TM) to achieve knowledge consolidation at both the instance and domain levels, enhancing anti-forgetting and generalization capacity.
Person re-identification (ReID) aims to match individuals across non-overlapping camera views. While existing ReID methods [1, 2, 3] achieve strong performance within fixed domains or short periods, they struggle to adapt to real-world scenarios [4, 5] where the domain continuously evolves. This limitation has motivated increasing research efforts toward lifelong person re-identification (LReID) [6, 7] that continuously acquire knowledge from sequentially arriving domains. The core challenge of LReID is to alleviate old knowledge forgetting while ensuring adaptation to new information.
Recently, LReID approaches have been focused on by many researchers to address this challenge, which can be categorized into two main branches: rehearsal-based [6, 8, 9] and rehearsal-free methods [7, 10, 11]. As depicted in Fig 1 (a), rehearsal-based schemes employ knowledge distillation to enforce representation alignment between the old and new model by jointly training old and new exemplars. They inevitably raise privacy concerns and consume computational resources [12, 13]. Alternatively, rehearsal-free solutions [14] adopt a distribution rehearser (DR) to model the last domain style for knowledge consolidation, limiting the model’s long-term anti-forgetting capacity, illustrated in Fig. 1 (b). Despite these advancements, two critical problems remain unresolved: 1) Specific Distribution Awareness. Significant distribution shifts caused by environmental variations (e.g., lighting conditions and scene backgrounds) lead to biased learning, where models exhibit strong performance on the current domain but severe degradation on historical ones. We believe that setting up a dedicated module with appropriate parameters [15] to retain the distribution of each domain without storing old exemplars is crucial for mitigating catastrophic forgetting. 2) Unified Knowledge Learning. Some methods [16, 17] align only the current representation with old representations derived from the old model for each instance. It is difficult to ensure that the old model learns a unified knowledge across previous domains to reduce inter-domain gaps. In general, specific distribution awareness and unified knowledge learning are highly significant for alleviating knowledge forgetting while ensuring effective adaptation to new information in the current domain.
To solve these issues, we propose a novel paradigm where domain-style modeling is performed to propagate domain-specific representations for knowledge unification and consolidation both at instance and domain levels with an exemplar-free and distillation-free framework, shown in Fig 1 (c). We aim to develop a model capable of transforming each instance of the current domain into domain-specific representations with the dedicated style of each domain. The unification of these domain-specific representations enables the model to better learn inter-domain distribution knowledge, improving its anti-forgetting and adaptation capacity.
Specifically, we propose a novel distribution-aware knowledge unification and association (DKUA) framework to explore domain-specific representations with the dedicated domain styles and the unified representation served as cross-domain shared knowledge for enhancing model’s anti-forgetting and adaptation capacity. Specifically, we design a distribution-aware model to transfer instance-level representations of the current domain into the domain-specific representations with the dedicated domain styles. Then, an adaptive knowledge consolidation (AKC) mechanism is proposed to dynamically unify domain-specific representations while ensuring consistency alignment between the current and all previous domains. To further alleviate catastrophic forgetting, we develop a unified knowledge association (UKA) mechanism, which explores the unified representation as a bridge to explicitly model inter-domain relationships, thereby reducing inter-domain gaps. Finally, we design distribution-based knowledge transfer (DKT) to prevent the current domain distribution from deviating from the unified distribution of all domains, improving anti-forgetting and adaptation capacity. Our contributions can be summarized as follows:
  • We propose a novel DKUA framework, a rehearsal-free and distillation-free approach that explores the association between domain-specific and unified representations for effective knowledge consolidation, significantly enhancing anti-forgetting and adaptation performance.
  • We propose an AKC module to adaptively unify domain-specific representations while preserving consistency between the current domain and all previous domains.
  • We propose a UKA mechanism, which utilizes the unified representation as a bridge to explicitly model inter-domain associations for reducing inter-domain gaps.
  • We propose a DKT mechanism to prevent the current domain distribution from deviating from the unified center, improving anti-forgetting and adaptation capacity.
II Related work
II-A Lifelong Person Re-Identification
Lifelong person re-identification (LReID) suffers from a key challenge, which involves a trade-off between the preservation of old knowledge and adaptation to new information. To solve this challenge, existing LReID approaches [18, 19, 20, 21] mainly consist of rehearsal-based and rehearsal-free schemes. Rehearsal-based approaches [8, 9] focus on knowledge distillation to achieve knowledge alignment by jointly training old and new exemplars. Despite their excellent performance in anti-forgetting ability, they inevitably raise privacy concerns and consume computational resources. Rehearsal-free solutions [7, 10] retain previous domain styles as old knowledge prior to preserve old knowledge forgetting, limiting the model’s long-term anti-forgetting capacity. In this paper, to overcome these limitation, we propose a rehearsal-free and distillation-free framework to explore each domain-specific style and the unified knowledge of all domains for enhancing anti-forgetting and generalization capacity.
II-B Distribution Awareness
In lifelong learning, distribution-aware methods [22] consolidate each domain-specific style to preserve the learned knowledge, rather than storing old exemplars. Some methods learn domain-specific knowledge by assigning the same structure, including the expert network [23, 24] and the adapter [11]. Additionally, some works [10] focus on modeling the data uncertainty to handle the out-of-distribution data for increasing the richness of data samples and alleviating old knowledge forgetting. In this paper, we investigate the dedicated style of each domain without storing historical exemplars. By transferring the current domain distribution to domain-specific styles, we repeatedly leverage old knowledge to alleviate catastrophic forgetting.
Figure 2: The pipeline of our DKUA framework. At the current training domain, the distribution-aware model transfers the robust representation of each instance to domain-specific representations $\{\theta^{i}\}_{i=1}^{t}$ with the dedicated domain styles. The AKC aims to learn the unified representation $\theta$ dynamically unify domain-specific representations while ensuring consistency alignment inter-domain. The UKA explores the unified representation as a bridge to explicitly model inter-domain associations $\{A_{u-i}\}_{i=1}^{t}$. The DKT prevents the current domain distribution from deviating from the unified distribution.
III Methodology
III-A Preliminary
In the rehearsal-free LReID task, a series of $T$ training domains $D=\{D^{t}\}_{t=1}^{T}$ is given to successively train the model. Specifically, $D^{t}=\{(x_{i},y_{i})\}_{i=1}^{N_{t}}$ consists of $N_{t}$ images $x_{i}$ and their identity labels $y_{i}$. At the $t$-th training domain, our model cannot access previous domains $\{D^{i}\}_{i=1}^{t-1}$ and only adopts the current domain $D^{t}$, including the Backbone $f_{B}$, a domain-style encoder (DSE) with tt-1 frozen transfer module (TM) $\{f_{M}^{i}\}_{i=1}^{t-1}$ and a current domain’s updatable transfer module (TM) $f_{M}^{t}$. To evaluate the model’s anti-forgetting and generalization ability, it is tested on a series of test domains $\{D_{test}^{i}\}_{i=1}^{T}$ from the Seen domain and Unseen domain $D_{un}$.
III-B Overview of DKUA
We propose a distribution-aware knowledge unification and association (DKUA) framework with a rehearsal-free and distillation-free framework, as illustrated in Fig. 2. Our approach consists of four key components: 1) distribution-aware model, 2) adaptive knowledge consolidation (AKC), 3) unified knowledge association (UKA), and 4) distribution-based knowledge transfer (DKT).
III-C Distribution-aware Model
To effectively preserve old knowledge, current methods rely on storing old exemplars, raising data privacy concerns. Unlike these approaches, we aim to learn the domain-specific instance-level representations with the dedicated domain style to preserve the obtained knowledge instead of storing old exemplars. Specifically, we propose a distribution-based model to transfer instance-level representations of the current domain into the domain-specific representations with the different domain styles for alleviating catastrophic forgetting. Distribution-aware model consists of the Backbone and a domain-style encoder (DSE), as illustrated in Fig. 2 (Distribution-aware Model).
Given a batch input image $x\in\mathbb{R}^{B\times C\times H\times W}$, the Backbone $f_{B}$ employs a vision transformer (ViT) with a depth setting of eleven to extract robust representations $f_{B}(x)\in\mathbb{R}^{B\times D}$ for each instance. [15] indicates that dynamic architectures can reduce catastrophic forgetting efficiently. Inspired by it, we design a domain-oriented encoder with $T$ transfer module (TM), which presents several advantages at the training domain $t$ as follows: 1) Current TM $f_{M}^{t}$ adaptively captures new knowledge from the current domain $t$. 2) Each previous TM $\{f_{M}^{i}\}_{i=1}^{t-1}$ with frozen parameters significantly retains the knowledge of each previous domain. 3) Each instance of the current domain is transformed into $t$ domain-specific representations with the dedicated domain styles, which preserves old knowledge without storing old exemplars. Domain-specific representations $\{\theta^{i}\}_{i=1}^{t}$ generated by domain-style encoder are calculated as:
$\theta^{i}=f_{M}^{i}(f_{B}(x))(i=1,2,\cdots,t)$ (1)
where $f_{M}^{i}$ indicates transfer module (TM) operations including a multi-head self-attention (MHSA), feed-forward network (FFN), and a linear layer.
III-D Adaptive Knowledge Consolidation
The domain-style encoder (DSE) captures domain-specific representations $\{\theta^{i}\}_{i=1}^{t}\in\mathbb{R}^{B\times D}$from the each instance of current domain $t$. To reduce inter-domain gaps, we aim to learn a unified representation by incorporating these domain-specific representations in the feature space. For this purpose, we propose an adaptive knowledge consolidation (AKC) mechanism to dynamically unify domain-specific representations while ensuring consistency alignment between the current and all previous domains, including knowledge unification (KU) and knowledge alignment (KA), as illustrated in Fig. 2 (Adaptive Knowledge Consolidation).
Knowledge Unification: We propose a knowledge unification (KU) to dynamic integrates domain-specific representations $\{\theta^{i}\}_{i=1}^{t}\in\mathbb{R}^{B\times D}$ with the dedicated domain style into a unified representation $\theta\in\mathbb{R}^{B\times D}$ served as cross-domain knowledge center. It effectively facilitates the learning of domain-invariant knowledge by the model to alleviate catastrophic forgetting. Specifically, we dynamically learn adaptive weighting coefficients $\omega\in\mathbb{R}^{B\times t}$ derived from robust representation $f_{B}(x)$ mappings, which can be formulated as:
$\omega=\delta(\phi(\mu(f_{B}(x))))$ (2)
where $\mu$($\cdot$) and $\phi$($\cdot$) are mean function and linear function, respectively. $\delta$($\cdot$) indicates softmax function. Then, we perform weighted linear combinations of domain-specific representations using dynamically learned weighting coefficients $\omega$ to obtain the unified representation $\theta$, expressed as:
$\theta=\sum_{i=1}^{t}\theta^{i}*\omega^{i}$ (3)
To ensure the new knowledge learning and the old knowledge anti-forgetting, we adopt the classical ReID loss including a cross-entropy loss $L_{\mathrm{CE}}$ and a triplet loss $L_{\mathrm{Tri}}$.
$L_{\mathrm{CE}}=-y\log(f_{c}(\theta))$ (4)
where $f_{c}$ is classifier and $y$ indicates identity label.
[…]
where $\theta_{p}$ and $\theta_{n}$ are the positive and negative samples related to an anchor sample $\theta$. $||\cdot||_{2}$ is LL2 norm. The classical ReID loss is formulated as:
$L_{\mathrm{ReID}}=L_{\mathrm{CE}}+L_{\mathrm{Tri}}$ (6)
Knowledge Alignment: Due to the significant inter-domain gaps between different training domains, directly integrating domain-specific representations into a unified representation through knowledge unification inevitably introduces bias. This bias shifts the unified representation away from the center of cross-domain knowledge. To address this, we propose a knowledge alignment (KA) mechanism to achieve knowledge consistency between the current domain and the previous domains, ensuring that the unified representation remains accurate and effective. At the training domain $t$, we constrain the current domain’s representation $\theta^{t}$ stepwise with the previous domain-specific representations $\{\theta^{i}\}_{i=1}^{t-1}$, ensuring that the unified representation remains centered across all domains. Specifically, we first calculate the similarity matrix $S_{t-(t-1)}$ between the current domain’s representation $\theta^{t}$ and the last domain’s representation $\theta^{t-1}$ using cosine similarity. Similarly, we derive the similarity matrix $S_{t-(t-2)}$ to $S_{t-1}$ for all preceding domains. Then, each similarity matrix $S_{t-(i-1)_{i=2}^{t-1}}$ is converted to a cosine distance metric $1-S_{t-(i-1)_{i=2}^{t}}$, enforcing distributional convergence between the current domain and each previous domain. This process is calculated as:
$S_{t-(i-1)}=Cos(\theta^{t},\theta^{t-1})$ (7)
$L_{\mathrm{KA}}=\frac{1}{t-1}\sum\limits_{i=2}^{t}\mu(1-S_{t-(t-1)})$ (8)
where $Cos(\cdot,\cdot)$ and $\mu$($\cdot$) are cosine similarity and mean function, respectively.
III-E Unified Knowledge Association
At the current training domain $t$, domain-specific representations and unified representation should meet the following two conditions: 1) Domain-specific representations $\{\theta^{i}\}_{i=1}^{t}$ should establish an affinity associations with the unified representation to measure similarity in different domain styles. 2) Unified representation $\theta$ serves as a bridge to bring different domain-specific representations $\{\theta^{i}\}_{i=1}^{t-1}$ closer, which improves domain-invariant representation learning. Thus, we propose a unified knowledge association (UKA) mechanism to explore cross-domain relationships, which aligns heterogeneous domain-specific representations and facilitates unified representation learning, as shown in Fig. 2 (Unified Knowledge Association). Specifically, we first calculate an association information $\{A_{u-i}\}_{i=1}^{t}$ between domain-specific representations and the unified representation. The association information estimates the relative distance between unified representation $\theta$ and domain-specific representations $\{\theta^{i}\}_{i=1}^{t}$, which can be formulated as:
$A_{u-i}=\delta(Cos(\theta^{t},\theta)/\lambda)(i=1,2,\cdots,t)$ (9)
where $\delta$ is the softmax function and $\lambda$ indicates the temperature parameter set to 0.1 [25]. To better learn a unified representation and reduce inter-domain gaps, we optimize the association information to minimize the distance between the current domain knowledge and the learned domain knowledge, calculated as follows:
$L_{\mathrm{UKA}}=\sum\limits_{i=2}^{t-1}KL(A_{u-t},A_{u-i})$ (10)
where $KL(\cdot,\cdot)$ is kullback leibler divergence [26].
III-F Distribution-based knowledge transfer
AKC and UKA strategies mainly focus on instance-level to promote the model’s anti-forgetting and generalization capability. To further boost model’s performance, we propose a distribution-based knowledge transfer (DKT) mechanism to prevent the current domain distribution from deviating from the unified distribution of all domains, improving anti-forgetting and adaptation capacity. Firstly, we calculate covariance matrix from all instance of each domain served as each domain distribution. Covariance matrix is formulated as:
$\Sigma^{t}=Cov(\theta_{a}^{t})$ (11)
where $Cov$ and $\theta_{a}^{t}\in\mathbb{R}^{N_{t}\times D}$ are covariance function and domain representations from all instances of current domain. Similarly, we derive covariance matrix of each domain $\{\Sigma^{i}\}_{i=1}^{t}$. Then, we transfer each domain distribution into a common covariance matrix $\Sigma^{1:t}$ as the unified distribution across all domains according to the following scheme:
$\Sigma^{1:t}=\Sigma^{1:t-1}\cdot\frac{|y^{1:t-1}|}{|y^{1:t}|}+\Sigma^{t}\cdot\frac{|y^{1:t}|-|y^{1:t-1}|}{|y^{1:t}|}$ (12)
where $y^{1:t-1}$ and $y^{1:t}$ indicate total number of classes in tt-1 domains and total number of classes in $t$ domains, respectively. Finally, we minimize the distribution distance between the unified distribution and the current domain distribution by kullback-leibler divergence, formulated as:
$L_{\mathrm{jdc}}=KL(\Sigma^{1:t},\Sigma^{t})$ (13)
The overall loss function $L$ of our method is formulated as:
$L=L_{\mathrm{ReID}}+L_{\mathrm{KA}}+L_{\mathrm{UKA}}+L_{\mathrm{DKT}}$ (14)
TABLE I: Dataset statistics of the LReID benchmark. Since the sampling procedure results in the numbers of train IDs being all 500, the original numbers of IDs are listed for comparison. ’-’ denotes that the dataset is not used for training
Type Datasets Scale Train IDs Test IDs
Seen Market-1501 [27] large 500(751) 750
CUHK-SYSU [28] mid 500(942) 2900
DukeMTMC [29] large 500(702) 1110
MSMT17_\_V2 [30] large 500(1041) 3060
CUHK03 [31] mid 500(700) 700
Unseen VIPeR [32] small
316
GRID [33] small
126
CUHK02 [34] mid
239
Occ_\_Duke [35] large
1100
Occ_\_REID [36] mid
200
PRID2011 [25] small
649
[…]
Method Type Market-1501 CUHK-SYSU DukeMTMC MSMT17_\_V2 CUHK03 Seen-Avg Unseen-Avg
mAP R@1 mAP R@1 mAP R@1 mAP R@1 mAP R@1 mAP R@1 mAP R@1
AKA[6] R-R 58.1 77.4 72.5 74.8 28.7 45.2 6.1 16.2 38.7 40.4 40.8 50.8 42.0 39.8
PTKP[37] R-R 64.4 82.8 79.8 81.9 45.6 63.4 10.4 25.9 42.5 42.9 48.5 59.4 51.2 49.1
PatchKD[16] R-R 68.5 85.7 75.6 78.6 33.8 50.4 6.5 17.0 34.1 36.8 43.7 53.7 45.1 43.3
KRKC[8] R-R 54.0 77.7 83.4 85.4 48.9 65.5 14.1 33.7 49.9 50.4 50.1 62.5 52.7 50.8
ConRFL[17] R-R 59.2 78.3 82.1 84.3 45.6 61.8 12.6 30.4 51.7 53.8 50.2 61.7 - -
LSTKC[38] R-R 54.7 76.0 81.1 83.4 49.4 66.2 20.0 43.2 44.7 46.5 50.0 63.1 51.3 48.9
C2R[39] R-R 69.0 86.8 76.7 79.5 33.2 48.6 6.6 17.4 35.6 36.2 44.2 53.7 - -
DRE[40] T-R 69.4 85.2 85.6 86.8 52.8 67.3 21.5 43.7 54.7 58.1 56.8 68.2 56.7 55.0
DKP[10] R-RF 60.3 80.6 83.6 85.4 51.6 68.4 19.7 41.8 43.6 44.2 51.8 64.1 49.9 46.4
DASK [14] R-RF 61.2 82.3 81.9 83.7 58.5 74.6 29.1 57.6 46.2 48.1 55.4 69.3 56.3 52.6
PAEMA[7] T-RF 71.7 86.9 90.0 91.3 66.0 79.0 31.4 55.3 50.0 50.8 61.8 72.7 61.1 57.9
DCR [24] T-RF 75.9 87.9 87.3 88.5 60.1 71.9 25.3 50.1 60.5 61.3 61.8 71.9 60.8 58.3
DKUA-ImageNet T-RF 80.3 89.7 87.5 88.6 58.7 73.4 27.0 50.6 63.1 64.8 63.3 73.4 62.9 60.5
DKUA T-RF 82.5 91.3 90.1 91.5 60.9 75.4 25.9 51.4 61.2 62.8 64.1 74.5 64.7 61.3
  • Type *-*, first * is Transformer-based (T) or ResNet-based (R) structure, second * indicates a Rehearsal-free (RF) or Rehearsal-based (R) scheme. The type T-RF is a transformer-based and rehearsal-free approach.
[…]
Method Type DukeMTMC MSMT17_\_V2 Market-1501 CUHK-SYSU CUHK03 Seen-Avg Unseen-Avg
mAP R@1 mAP R@1 mAP R@1 mAP R@1 mAP R@1 mAP R@1 mAP R@1
AKA[6] R-R 42.2 60.1 5.4 15.1 37.2 59.8 71.2 73.9 36.9 37.9 38.6 49.4 41.3 39.0
PTKP[37] R-R 54.8 70.2 10.3 23.3 59.4 79.6 80.9 82.8 41.6 42.9 49.4 59.8 50.8 48.2
PatchKD[16] R-R 58.3 74.1 6.4 17.4 43.2 67.4 74.5 76.9 33.7 34.8 43.2 54.1 44.8 43.3
KRKC[8] R-R 50.6 65.6 13.6 27.4 56.2 77.4 83.5 85.9 46.7 46.6 50.1 61.0 52.1 47.7
ConRFL[17] R-R 34.4 51.3 7.6 20.1 61.6 80.4 82.8 85.1 49.0 50.1 47.1 57.4 - -
LSTKC[38] R-R 49.9 67.6 14.6 34.0 55.1 76.7 82.3 83.8 46.3 48.1 49.6 62.1 51.7 49.5
C2R[39] R-R 59.7 75.0 7.3 19.2 42.4 66.5 76.0 77.8 37.8 39.3 44.7 55.6 - -
DRE[40] T-R 59.7 74.2 18.7 34.8 65.4 82.7 84.8 86.7 51.9 53.2 56.1 66.3 57.1 54.9
DKP[10] R-RF 53.4 70.5 14.5 33.3 60.6 81.0 83.0 84.9 45.0 46.1 51.3 63.2 51.3 47.8
DASK [14] R-RF 55.7 74.4 25.2 51.9 71.6 87.7 84.8 86.2 48.4 49.8 57.1 70.0 57.7 53.9
PAEMA[7] T-RF 67.2 79.8 26.0 49.4 69.8 85.8 89.9 91.0 49.3 49.7 60.4 71.1 60.5 57.6
DCR [24] T-RF 64.1 77.2 25.4 44.9 70.6 84.5 86.1 88.2 54.2 58.7 60.1 70.7 61.6 59.2
DKUA-ImageNet T-RF 66.1 79.2 29.6 54.6 68.5 83.6 86.4 87.9 61.8 63.4 62.5 73.7 64.2 61.4
Ours T-RF 69.5 81.8 31.8 57.6 73.7 87.9 87.6 88.1 59.5 60.9 64.4 75.3 65.7 62.6
  • Type *-*, first * is Transformer-based (T) or ResNet-based (R) structure, second * indicates a Rehearsal-free (RF) or Rehearsal-based (R) scheme. The type T-RF is a transformer-based and rehearsal-free approach.
IV Experiments
IV-A Experiments Setting
Datasets: All experiments are conducted on a challenging benchmark [6] consisting of a Seen domain and an Unseen domain. Seen domain includes Market1501[27], CUHK-SYSU [28], DukeMTMC [29], MSMT17_\_V2 [30] and CUHK03 [31]. Further, unseen domain consists of VIPeR [32], GRID [33], CUHK02 [34], Occ_\_Duke [35], Occ_\_REID [36], and PRID2011 [25]. Following [6], two classic training orders are used to imitate varying domain gaps and assess the model’s performance. More detailed statistics for Seen domain and an Unseen domain are provided in Table I.
[…]
Evaluation Metrics. To measure performance in each domain, we adopt two standard metrics: Mean Average Precision (mAP) [27] and Rank-1 (R@1) accuracy [44]. Additionally, to evaluate the model’s lifelong learning ability, we average the mAP and R@1 scores over both seen and unseen domains, providing a holistic assessment of its generalization.
Figure 3: Anti-forgetting curves on Market-1501.
IV-B Compared Methods
We compare the proposed DKUA with two streams of methods: rehearsal-based and rehearsal-free methods. Rehearsal-based approaches consist of AKA[6], PTKP[37], PatchKD[16], KRKC[8], ConRFL[17], LSTKC[38], C2R[39], and Vit-based DRE[40]. Rehearsal-free approaches include ResNet-based DKP[10] and DASK[14], Vit-based PAEMA[7] and DCR [24]. Experimental results for training order-1 and training order-2 are presented in Table II and Table III, respectively. DKUA-ImageNet employ Vit-B/16 [41], initializing the model weights with pre-training on ImageNet-21K [45].
IV-C Performance Assessment on Seen domain
Compared to rehearsal-based methods: Rehearsal-based methods employ knowledge distillation to align knowledge by jointly training on both old and new exemplars. As shown in Table II and Table III, achieves superior performance compared to ResNet-based methods across all five training domains, with Seen-Avg gains of 13.9% mAP/11.4% R@1, and 14.3% mAP/13.2% R@1 under the two training orders, respectively. Furthermore, compared to the Vit-based DRE approach. our DAFC outperforms it by 7.3% mAP/6.3% R@1, and 8.3% mAP/9.0% R@1 in Seen-Avg across the two training orders. These improvements arise from our distribution-aware model design, which effectively captures domain-specific styles without relying on stored old exemplars, achieving an optimal trade-off between anti-forgetting and adaptive learning.
Compared to rehearsal-free methods: Rehearsal-free solutions retain previous domain styles as old knowledge to preserve old knowledge. As shown in Table II and Table III, compared to ResNet-based DKP and DASK, our DKUA achieves significant gains of 8.7% mAP/5.2% R@1 and 7.3% mAP/5.3% R@1 in Seen-Avg under two training orders, respectively. Specifically, DKUA outperforms DASK on four out of five datasets in both mAP and R@1 within the Seen domain. However, DASK outperforms our DKUA on MSMT17_V2, as it employs a distribution-rehearsing mechanism that effectively models MSMT17_V2’s distribution to achieve consistency alignment by knowledge distillation, enhancing the model’s learning capacity on a fixed domain. Compared with Vit-based PAEMA and DCR methods, our DKUA yields Seen-Avg improvements of 2.3% mAP/1.8% R@1 and 4.0% mAP/4.2% R@1 under the two training orders. These results highlight the advantage of our distribution-aware design, which captures diverse domain styles to improve cross-domain adaptability. By establishing affinity associations between unified and domain-specific knowledge at both instance and domain levels, our model effectively reduces inter-domain gaps while balancing anti-forgetting and adaptation.
Anti-forgetting Curves: To evaluate the anti-forgetting capability, we perform experiments on Market-1501 under the training order-1, as shown in Fig. 3. These results demonstrate that compared methods exhibit significant performance degradation after training on the large-scale MSMT17 dataset (Training step=4). In contrast, our DKUA maintains relatively stable mAP and R@1 performance as the training step increases. This significant achievement demonstrates that our DKUA effectively captures domain-specific knowledge for preserving old domain distributions while facilitating unified knowledge learning across all domains, significantly mitigating catastrophic forgetting.
Figure 4: Generalization tendency on Unseen domain.
IV-D Performance Assessment on Unseen domain
Compared to rehearsal-based methods: As shown in the Unseen-Avg in Table II and Table III, our DKUA significantly outperforms rehearsal-based methods on the Unseen domain, with an Unseen-Avg incremental gain of 12.0% mAP/10.5% R@1, and 12.4% mAP/13.1% R@1 improvement on two training orders. We consider that limiting the performance of the ResNent-based model presents two aspects: 1) The representation ability of the model is poor, making it difficult to distinguish people with similar identities. 2) The model insufficiently consolidates cross-domain unified knowledge, resulting in a decrease in generalization performance on the Unseen domain. Furthermore, compared to the ViT-based DRE method, our DAFC achieves superior Seen-Avg performance by 7.3% mAP/6.3% R@1 and 8.3% mAP/9.0% R@1 across the two training orders. In general, our adaptive knowledge consolidation (AKC) design effectively explores domain-specific representations for each instance to mine the unified representation across all domains, improving the model’s performance on the unseen domain.
Compared to rehearsal-free methods: In Table II and Table III, our DKUA achieves significant performance, obtaining 3.6%/3.0% and 4.1%/3.4% Seen-Avg mAP/R@1 improvement on two training orders compared to the Vit-based PAEMA. This can be attributed to our distribution-aware model design, which maintains learned domain-specific styles at each training step. By transforming instances of the current domain into different domain styles, it guides the model to learn both cross-domain unified knowledge and identity discriminative information, significantly enhancing the model’s generalization capability.
Generalization of Unseen Domain: As illustrated in Fig. 4, compared to existing methods, our DKUA method achieves superior performance growth on the Unseen domain across the training steps without using a knowledge distillation strategy or storing old exemplars. In summary, our DKUA significantly achieve knowledge consistency alignment between domain-specific knowledge and the unified knowledge, facilitating the model to capture domain invariant knowledge for improving generalization capability.
TABLE IV: Ablation studies of different components on training order-1.
Baseline AKC UKA DKT Seen_\_Avg Unseen_\_Avg
mAP R-1 mAP R-1
$\surd$ 58.2 65.8 57.5 54.6
$\surd$ $\surd$ 59.5 67.3 58.9 55.8
$\surd$ $\surd$ $\surd$ 62.6 72.1 62.4 59.7
$\surd$ $\surd$ $\surd$ $\surd$ 64.1 74.5 64.7 61.3
Figure 5: The t-SNE visualization results of the unified representation across five domains.
Figure 6: Visualization of retrieval results. The left column displays the query image under challenging conditions, such as low-light, occlusion, back views, and side views. In the retrieved results, correctly matched pedestrians are marked with green boxes, while incorrect matches are highlighted in red box. Compared to the DASK method, our DKUA retrieves more pedestrians with the same identity, demonstrating superior performance.
IV-E Ablation Studies
To assess the contributions of individual components in our DKUA, we conduct ablation studies to determine the impact of Baseline, adaptive knowledge consolidation (AKC), unified knowledge association (UKA), distribution-based knowledge transfer (DKT) on model performance. Baseline consists of distribution-aware model and knowledge unification. These experiments are performed on unseen and seen domains under training order-1, as illustrated in Table IV.
Performance of Different Components: In Table IV, the Baseline demonstrates that our distribution-aware model and knowledge unification strategy effectively capture domain-specific distribution and consolidate unified representation across all domains for alleviating catastrophic forgetting. Compared to the second row, our AKC enforces narrowing the distance between domains, ensuring that unified representation learning remains accurate and effective. Compared to the second and third row, our unified knowledge association mechanism achieves a significant improvement of 3.1% mAP/4.8% R@1, and 3.5% mAP/4.1% R@1. This demonstrates that our UKA effectively associates knowledge between domain-specific and unified representations based on each instance to reduce inter-domain gaps for enhancing the model’s anti-forgetting and generalization. Compared to the third and fourth rows, our DKT design enables the model to transfer domain-specific distributions into unified knowledge based on all exemplars of each domain. Each independent component in our DKUA plays a crucial role. When the four components are merged, our DKUA significantly improves both the anti-forgetting and generalization capacity of the model.
Performance of unified representation: To validate the unified representation learning ability of our method, we perform the t-SNE visualization results of PTKP, KRKC, DKP, and our DKUA across five domains, as shown in Fig. 5. The KRKC and DKP methods struggle to accumulate sufficient cross-domain unified knowledge, which limits their anti-forgetting capacity. The PTKP method inadequately separates identity information within the domain, causing a decrease in matching accuracy. Compared to these methods, our DKUA presents two advantages: 1) Full interaction of cross-domain information demonstrates that our DKUA significantly exploits the unified representation across all domains for reducing inter-domain gaps. 2) Differentiated identity information reveals that our method dynamically consolidates the unified representation that can effectively distinguish similar identities across all domains. In summary, our DKUA effectively consolidates the unified representation with the characteristics of cross-domain sharing and identity discrimination to enhance the model’s anti-forgetting and matching accuracy performance.
TABLE V: Experimental results regarding training time, Parameters, and FLOPs.
Method Training time (h-min) Parameters(M) FLOPs(G)
DRE[40] 20h-18min 99.83 21.2
PAEMA[7] 6h-20min 86.03 11.3
DCR [24] 19h-10min 129.5 12.6
DKUA-ImageNet 6h-35min 102.7 18.3
DKUA 3h-22min 39.67 8.1
IV-F Analysis of Computational Complexity
Complexity Analysis of transfer module (TM): Each domain is equipped with a dedicated Transfer Module (TM) to capture its domain-specific distribution, thereby mitigating catastrophic forgetting. To quantify the increase in model complexity, we compute the FLOPs and parameters of each TM. The TM adopts the structure of a final transformer layer, which includes multi-head self-attention (MHSA), a feed-forward network (FFN), residual connections, and layer normalization. A single TM introduces 228.7 M FLOPs and 1.774 M parameters, which uses a small amount of memory compared to storing old exemplars [40]. This design allows our DKUA model to effectively learn and retain domain-specific knowledge, successfully alleviating catastrophic forgetting.
Efficiency of Computational Complexity: We conduct experiments to evaluate computational complexity in comparison with ViT-based methods, including DRE [40], PAEMA [7], DCR [24], and DKUA-ImageNet, as summarized in Table V. Training time is reported in hours (h) and minutes (min). Although our DKUA-ImageNet produces higher FLOPs and parameter counts than DRE, DCR, and PAEMA, it achieves a shorter training time than DRE and DCR. More importantly, DKUA-ImageNet delivers significantly better performance, with improvements of at least 1.5% mAP/0.7% R-1 and 2.1% mAP/2.6% R-1 on Seen-Avg, and at least 1.8% mAP/2.2% R-1 and 2.6% mAP/2.2% R-1 on Unseen-Avg for training order-1 and order-2, as shown in Tables II and III. To reduce the model’s parameters and FLOPs, our DKUA introduces a ViT-Small/16 [41] architecture as the backbone. Our strategies (such as AKC, UKA, and DKT) with a rehearsal-free and distillation-free scheme are well-suited to the backbone, significantly enhancing the model’s anti-forgetting and generalization abilities in Tables II and III.
IV-G Visualization
To further validate the effectiveness of our DKUA, we conduct a qualitative comparison with the DASK method, as shown in Fig 6. The left column displays challenging query images, including cases with low-light conditions, occlusion, back and side views. In Fig. 6, correctly matched pedestrians are marked with green boxes, while incorrect matches are highlighted in red. Under low-light conditions (first row), our DKUA accurately retrieves eight out of ten individuals with the same identity, despite the limited visual information available. In the case of occlusion (second and third rows), the DASK method produces at least four incorrect matches out of ten individuals, whereas DKUA correctly retrieves nine. For back and side views (fourth and fifth rows), our method successfully handles the perspective change issue and outperforms DASK in identifying the same identity. In the side-view scenario with occlusion (sixth row), DASK correctly matches only four pedestrians, while DKUA achieves perfect retrieval (10/10). These results demonstrate that DKUA achieves superior retrieval performance by consistently identifying more pedestrians of the same identity across diverse challenging conditions. This advantage stems from its ability to effectively learn a unified representation that captures cross-domain shared features while maintaining identity discriminability, thereby enhancing the model’s anti-forgetting capability and matching accuracy.
V Conclusions
We propose a novel distribution-aware knowledge unification and association (DKUA) framework with an exemplar-free and distillation-free framework. To consolidate different domain distributions into unified knowledge, we develop a distribution-aware model and an adaptive knowledge consolidation module to capture domain-specific representation based on the instance level and adaptively transfer different domain knowledge into the unified representation. Meanwhile, the unified knowledge association and distribution-based knowledge transfer mechanisms are proposed to unite domain-specific knowledge to the unified representation through instance-wise correlation and domain-wise distribution alignment. Experiments show our DKUA outperforms state-of-the-art LReID methods under two training orders and achieves strong generalization on the Unseen domain.